A car is driving east for 50km, North for 30km, then 30
1
Expert's answer
2021-08-02T08:40:16-0400
"x=50+25\sin{30}=62.5\ km\\y=30+25\cos{30}=51.65\ km\\
s=\sqrt{62.5^2+51.65^2}=81.1\ km\\"